OpenStudy (anonymous):

Which one is a rule for the nth term of an arithmetic sequence with a8 = 56 and d = 5?

OpenStudy (anonymous):

A. an = 16 − 5n B. an = 26 + 5n C. an = 26 − 5n D. an = 16 + 5n